# Large Language Model Inference Acceleration: A Comprehensive Hardware Perspective

### v2

@@ 3.1.3 GPU
Weight-Activation Quantization. In addition to hardware units that support FP16 computations, NVIDIA GPUs also provide hardware units that support INT4, INT8, and FP8 computations. The number of these computation units can be 2$\times$ and 4$\times$ greater than FP16 on each chip. Compared to weight-only quantization, weight-activation quantization can utilize INT4, INT8, and FP8 computations, thereby maximizing the peak computational performance of the GPU. Since the prefill phase in LLM inference is compute-bound, weight-activation quantization can significantly enhance performance during this stage. LLM.int8 [111] uses vector-wise quantization with separate normalization constants for each inner product in the matrix multiplication, to quantize most of the features. For the outliers, it isolates the outlier feature dimensions into a 16-bit matrix multiplication while still more than 99.9% of values are multiplied in 8-bit. For BLOOM-176B model, LLM.int8 achieves 4.05 tokens/s, 30.3 tokens/s and 109.77 tokens/s for batch size 1, 8 and 32, respectively, on 3 A100 GPUs in decode phase. The inference speed is slightly slower but close to 16-bit inference with less GPU consumption. SmoothQuant [112] enables 8-bit weight and 8-bit activation (W8A8) quantization for LLMs. Based on the fact that weights are easy to quantize while activations are not, SmoothQuant smooths the activation outliers by offline migrating the quantization difficulty from activations to weights with a mathematically equivalent transformation. SmoothQuant enables an INT8 quantization of both weights and activations for all the matrix multiplications in LLMs. SmoothQuant achieves up to 1.56$\times$ speedup and 2$\times$ memory reduction for LLMs with negligible loss in accuracy. QUIK [113] is for the first time, that the majority of inference computations for LLMs can be performed with both weights and activations being cast to 4 bits. QUIK compresses most of the weights and activations to 4-bit, while keeping some outlier weights and activations in higher-precision. It also provides GPU kernels matching the QUIK format with highly-efficient layer-wise runtimes, which lead to practical end-to-end throughput improvements of up to 3.4$\times$ relative to FP16 execution in prefill phase.
@@
Prevalent quantization schemes (e.g., W8A8) cannot fully leverage the capabilities of modern GPUs, such as 4-bit integer operators, resulting in sub-optimal performance. To maximize the throughput, Atom [114] significantly boosts serving throughput by using low-bit operators and considerably reduces memory consumption via low-bit quantization. It attains high accuracy by applying a novel mixed-precision and fine-grained quantization process. For single batch inference, Atom can achieve about 30 tokens/s for on a NVIDIA RTX 4090 GPU. Atom improves end-to-end throughput by up to 7.73$\times$ compared to the FP16 and by 2.53$\times$ compared to INT8 quantization, while maintaining the same latency target.
@@
Compared to integer quantization, floating-point (FP) quantization can better handle long-tail or bell-shaped distributions, and it has emerged as a default choice in many hardware platforms. LLM-FP4 [115] quantizes both weights and activations in LLMs down to 4-bit floating-point values (W4A4) with negligible accuracy loss. Due to the lack of PF4 computing unit in GPUs, its decoding speed maybe slower than FP16 baseline.
3.1.4 FPGA
@@
Weight-Activation Quantization. FlexRun [116] uses 8-bit quantization (W8A8), conducts an in-depth design space exploration to find the best accelerator architecture for a target LLM model, and automatically reconfigures the accelerator based on the exploration results. With the implementation on Intel Stratix 10 GX and MX FPGAs, FlexRun outperforms the current state-of-the-art FPGA-based accelerator by 1.15$\times$–1.50$\times$ for GPT2, respectively. Compared to Nvidia’s V100 GPU, FlexRun achieves 2.69$\times$ higher performance on average for various GPT2 models. HLSTransform [117] uses HLS to design a FPGA accelerator and synthesis combined with pipelining, memory unrolling, and memory partitioning and transfer optimizations, with the addition of 8-bit integer quantization (W8A8). On a tiny model with 110 million parameters, HLSTransform achieves 57.11 tokens/s on Xilinx Virtex UltraScale+ VU9P FPGA. SECDA-LLM [118] utilizes quantization (W3A8) and designs an efficient FPGA-based LLM accelerators for the llama.cpp inference framework. By deploying on the PYNQ-Z1 board, it achieves 0.588 tokens/s for the TinyLlama model (1.1B). Chen et al. [119] investigate the feasibility and potential of model-speciic spatial acceleration for LLM inference on FPGAs. They introduce a comprehensive analytical model to estimate the LLM inference performance of FPGA accelerator with W4A8 quantization, and provide a library of high-level synthesis (HLS) kernels that are composable and reusable. For Llama2-7B, during prefilling phase, they can achieves about 213 tokens/s, 43 tokens/s, and 320 tokens/s on Xilinx Alveo U280, VCK5000, and VHK158 FPGAs, respectively. During decode stage, they can achieves about 200 tokens/s, 40 tokens/s, and 333 tokens/s on Xilinx Alveo U280, VCK5000, and VHK158 FPGAs, respectively.
3.1.5 ASIC
@@
Weight-Only Quantization. Despite the memory footprint reduction achieved by weight-only quantization, the actual computing performance is not really improved due to dequantization from integer to float. FIGNA [122] proposes dedicated FP-INT arithmetic units designed specifically for FP-INT MAC operations and integrates them on the accelerator. FIGNA with FP16-INT4 provides 3.2768 TOPS computing power and 26.58W power consumption by considering all memory access in 28nm at 100MHz. Estimated result shows that for OPT-6.7B it can achieve 21.332 tokens/s in decode stage. Different from normal quantization methods, MECLA [123] proposes a parameter-efficient scaling sub-matrix partition method (SSMP) to decompose large weight matrices into several tiny-scale source sub-matrices (SS) and derived sub-matrices (DS). For memory issues, SSMP avoids accessing the full weight matrix but only requires small SS and DS scaling scalars. For computation issues, the proposed accelerator fully exploits the intermediate data reuse of matrix multiplication via on-chip matrix regrouping, inner-product multiplication re-association, and outer-product partial sum reuse. Totally, it can reduce 83.6% memory access and 72.2% computation. MECLA provides 14.008 TOPS computing power and $\sim$96W (1.9763W+94W) power consumption by considering all memory access under 28nm. For Llama2-7B and BLOOM-7B, compared to NVIDIA V100 GPU, MECLA achieves 6.74$\times$ and 5.91$\times$ inference speedup ($\sim$161 tokens/s and 141 tokens/s, respectively).
@@
Weight-Activation Quantization. Based on the key insight that outliers are important while the normal values next to them are not, OliVe [124] adopts an outlier-victim pair (OVP) quantization and handles outlier values locally with low hardware overheads. This enables a memory-aligned W4A4/W8A8 quantization, which can be efficiently integrated to the existing hardware accelerators like systolic array and tensor core. OliVe provides 0.71 TOPS computing power and $\sim$8W (0.2806W+7.9872W) power consumption by considering all memory access under 22nm. Estimated results shows that for OPT-6.7B it can achieve 9.173 tokens/s in decode stage. Li et al. [125] uniformly group weights and activations to ensure workload balance for hardware, and propose two approaches called channel sorting and channel selection to enhance the performance of quantization. It provides 1.43 TOPS computing power and $\sim$8.5W (0.472W+7.9872W) power consumption by considering all memory access under 65nm. Estimated results shows that for OPT-6.7B it can achieve 19.733 tokens/s in decode stage. Tender [126] decomposes weight and activation matrices by groups with different size to smooth the impact of outliers. And the format of scale factors are powers of two apart, which avoids explicit dequantization and extension to the commodity tensor compute hardware. It is 7.174W (1.60W+5.574W) power consumption by considering HBM2 memory access under 28nm. Result shows that for OPT-6.7B, Tender achieves 1.33$\times$ speedup (53.33 tokens/s) than NVIDIA A100 GPU.
3.1.6 PIM/NDP
@@
Figure 4: LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with quantization.
3.1.7 Quantitative Comparison
We first compare the power consumption, inference speed and energy efficiency for different hardware platforms, in Figure 4. For quantization, power consumption ranges from 3W to 450W, with inference speeds between 3 tokens/s and 1998 tokens/s. The energy efficiency ranges from 0.0167 tokens/J to 46.66 tokens/J. T-MAC [98] (CPU) achieves the lowest power consumption with 3W and Guo et al. [127] (PIM/NDP) achieves the highest throughput with batch size 1 (pre-silicon simulation result). And Guo et al. [127] also achieves the highest energy efficiency with 46.66 tokens/J.
@@ 3.2.1 Overview
Figure 5: Sparsity and sparsity patterns.
Table 4: Sparsity on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@ 3.2.2 CPU
Activation Sparsity. Activation sparsity is determined by activation functions. Commonly using SwiGLU [215] and GeGLU [216] exhibits limited sparsity for LLMs, but simply replacing these functions with ReLU fails to achieve sufficient sparsity. Turbo Sparse [130] proposes the dReLU activation function to improve LLM activation sparsity, along with a high-quality training data mixture ratio to facilitate effective sparsity. By applying their sparsity method to the Mistral and Mixtral models, only 2.5 billion (35.7%) and 4.3 billion (9.2%) parameters are activated per inference iteration, respectively. For Mistral-7B, Turbo Sparse achieves 8.71 tokens/s and 9.94 tokens/s on Intel i9-14900HX processor and Intel i7-12700K processor, respectively. For Mixtral-47B with 4-bit quantization, Turbo Sparse achieves 16.1 tokens/s, 11.98 tokens/s and 11.1 tokens/s on Intel i9-14900HX, Intel i7-12700K and SnapDragon 8 Gen3, respectively. ProSparse [131] also introduces activation function substitution, progressive sparsity regularization, and activation threshold shifting to help non-ReLU LLMs obtain high activation sparsity without performance degradation. For Llama2-7B and Llama2-13B, ProSparse achieves high sparsity of 89.32% and 88.80%, and 16.3 tokens/s and 8.67 tokens/s, respectively, based on PowerInfer [195] framework.
3.2.3 GPU
@@
Weight Sparsity. LLM-pruner [132], adopts structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the majority of the LLM’s functionality. To this end, the performance of pruned models can be efficiently recovered through tuning techniques, LoRA, in merely 3 hours, requiring only 50K data. We validate the LLM-Pruner on three LLMs, including Llama, Vicuna, and ChatGLM, and demonstrate that the compressed models still exhibit satisfactory capabilities in zero-shot classification and generation.
@@
LLMs can be pruned to at least 50% sparsity in one-shot, without any retraining, at minimal loss of accuracy. SparseGPT [133] requires a sophisticated weight update procedure in an iterative pruning process. Wanda [134] prunes weights with the smallest magnitudes multiplied by the corresponding input activations, on a per-output basis. Notably, Wanda requires no retraining or weight update, where pruning process is faster. Besides unstructured pattern, these two methods generalizes to semi-structured N:M (2:4 and 4:8) patterns. E-Sparse [135] introduces entropy to quantify the information richness within each channel (intra-channel) of the input features, and adopts it to enhance the feature norms (crosschannel) as a metric to evaluate parameter importance. Furthermore, it proposes Channel Shuffle to reorder the information distribution in LLMs to obtain N:M Sparsity with less information loss. 2:4 sparsity as supported by NVIDIA GPUs of generation Ampere and newer theoretically offers 2$\times$ acceleration of matrix multiplications. In practical, 2:4 sparsity can achieve 1.54$\times$-1.79$\times$ speedup for MatMul, and end-to-end speedups are about 1.21$\times$-1.25$\times$ (due to some extra overheads from e.g. attention).
@@
Based on the key observation that the bottleneck of LLM inference is the skinny matrix multiplications, Flash-LLM [136] proposes a general Load-as-Sparse and Compute-as-Dense methodology for unstructured sparse matrix multiplication. Flash-LLM proposes a new sparse format called Tiled-CSL to relieve the memory bandwidth bottleneck and support the tile-by-tile SpMM execution with tensor cores. For OPT-30B, Flash-LLM achieves 80% sparsity with 1.44% accuracy decrease and about 290 tokens/s, 500 tokens/s, 800 tokens/s, and 1187 tokens/s on single A100 GPU with batch sizes 8, 16, 32, and 64, respectively.
@@
Agarwalla et al. [137] combine the SparseGPT one-shot pruning method and sparse pretraining to pretrain a high sparsity LLM. They deploy model on GPU and CPU by utilizing Neural Magic’s DeepSparse engine and Neural Magic’s nm-vllm engine, respectively. For Llama-7B, on NVIDIA A10 GPU, they achieve 44.4 tokens/s and 47.9 tokens/s with 50% sparsity and 70% sparsity, respectively. On AMD EPYC 9R14 Processor, they achieve 4.4 tokens/s and 6.9 tokens/s with 50% sparsity and 70% sparsity, respectively.
@@
Existing methods require costly retraining, forgo LLM’s in-context learning ability, or do not yield wall-clock time speedup on modern hardware. DejaVu [138] predicts contextual sparsity on the fly given inputs to each layer, along with an asynchronous and hardware-aware implementation that speeds up LLM inference. For OPT-175B model, DejaVu achieves up to 75% sparsity and 50 tokens/s on 8 A100-80GB GPUs with batch size 1, which is over 2$\times$ and 6$\times$ faster than FasterTransformer and Hugging Face implementation, respectively.
@@
Attention Sparsity. During the prefilling phase of LLM inference, attention computation complexity scales quadratically with input sequence length. Given limited GPU computing and memory resources, attention sparsification can reduce the number of attention values to accelerate prefilling phase. For static sparsity, Sparse Transformer [139], StreamingLLM [141], Bigbird [140], and Longformer [142] use the manual combination of global and local patterns to replace the full attention patterns. The local pattern captures the local context of each token within a fixed size or stride while the global pattern captures the relationship between the specific tokens to all other tokens. For Llama2-7B and Llama2-13B models, StreamingLLM achieves 15.38-32.26 tokens/s and 9.43-20.83 tokens/s on single NVIDIA A6000 GPU, respectively. For dynaimic sparsity, Adaptively Sparse Attention [143] replaces softmax with $\alpha$-entmax, a differentiable generalization of softmax that allows low-scoring words to receive precisely zero weight and drops parts of the context that are no longer required for future generation. Reformer [144] replaces dot-product attention by using locality-sensitive hashing, changing the complexity from $O(L^{2})$ to $O(LlogL)$, where $L$ is the sequence length. Sparse Flash Attention [145] extends FlashAttention [217] GPU kernel and encompasses key/query dropping and hashing-based attention. Sparse Sinkhorn Attention [146] adopts a learned sorting network to align keys with their relevant query buckets, ensuring that attention is computed only between the corresponding query-key pairs. H2[147] observes that a small portion of tokens (called Heavy Hitters, H2) contributes most of the value when computing attention scores. H2O introduces a dynamic attention sparsification method to adopt KV cache eviction policy that dynamically retains a balance of recent and H2 tokens. For OPT-6.7B model, H2O with 20% H2 achieves 30.4 tokens/s on single NVIDIA T4 GPU.
3.2.4 FPGA

### v3

@@ 3.1.3. GPU
Weight-Activation Quantization. In addition to hardware units that support FP16 computations, NVIDIA GPUs also provide hardware units that support INT4, INT8, and FP8 computations. The number of these computation units can be 2$\times$ and 4$\times$ greater than FP16 on each chip. Compared to weight-only quantization, weight-activation quantization can utilize INT4, INT8, and FP8 computations, thereby maximizing the peak computational performance of the GPU. Since the prefill phase in LLM inference is compute-bound, weight-activation quantization can significantly enhance performance during this stage. LLM.int8 (dettmers2022gpt3, ) uses vector-wise quantization with separate normalization constants for each inner product in the matrix multiplication, to quantize most of the features. For the outliers, it isolates the outlier feature dimensions into a 16-bit matrix multiplication while still more than 99.9% of values are multiplied in 8-bit. For BLOOM-176B model, LLM.int8 achieves 4.05 tokens/s, 30.3 tokens/s and 109.77 tokens/s for batch size 1, 8 and 32, respectively, on 3 A100 GPUs in decode phase. The inference speed is slightly slower but close to 16-bit inference with less GPU consumption. SmoothQuant (xiao2023smoothquant, ) enables 8-bit weight and 8-bit activation (W8A8) quantization for LLMs. Based on the fact that weights are easy to quantize while activations are not, SmoothQuant smooths the activation outliers by offline migrating the quantization difficulty from activations to weights with a mathematically equivalent transformation. SmoothQuant enables an INT8 quantization of both weights and activations for all the matrix multiplications in LLMs. SmoothQuant achieves up to 1.56$\times$ speedup and 2$\times$ memory reduction for LLMs with negligible loss in accuracy. QUIK (ashkboos2023towards, ) is for the first time, that the majority of inference computations for LLMs can be performed with both weights and activations being cast to 4 bits. QUIK compresses most of the weights and activations to 4-bit, while keeping some outlier weights and activations in higher-precision. It also provides GPU kernels matching the QUIK format with highly-efficient layer-wise runtimes, which lead to practical end-to-end throughput improvements of up to 3.4$\times$ relative to FP16 execution in prefill phase.
@@
Prevalent quantization schemes (e.g., W8A8) cannot fully leverage the capabilities of modern GPUs, such as 4-bit integer operators, resulting in sub-optimal performance. To maximize the throughput, Atom (zhao2024atom, ) significantly boosts serving throughput by using low-bit operators and considerably reduces memory consumption via low-bit quantization. It attains high accuracy by applying a novel mixed-precision and fine-grained quantization process. For single batch inference, Atom can achieve about 30 tokens/s for on a NVIDIA RTX 4090 GPU. Atom improves end-to-end throughput by up to 7.73$\times$ compared to the FP16 and by 2.53$\times$ compared to INT8 quantization, while maintaining the same latency target.
@@
Compared to integer quantization, floating-point (FP) quantization can better handle long-tail or bell-shaped distributions, and it has emerged as a default choice in many hardware platforms. LLM-FP4 (liu2023llm, ) quantizes both weights and activations in LLMs down to 4-bit floating-point values (W4A4) with negligible accuracy loss. Due to the lack of PF4 computing unit in GPUs, its decoding speed maybe slower than FP16 baseline.
3.1.4. FPGA
@@
Weight-Activation Quantization. FlexRun (hur2023fast, ) uses 8-bit quantization (W8A8), conducts an in-depth design space exploration to find the best accelerator architecture for a target LLM model, and automatically reconfigures the accelerator based on the exploration results. With the implementation on Intel Stratix 10 GX and MX FPGAs, FlexRun outperforms the current state-of-the-art FPGA-based accelerator by 1.15$\times$–1.50$\times$ for GPT2, respectively. Compared to Nvidia’s V100 GPU, FlexRun achieves 2.69$\times$ higher performance on average for various GPT2 models. HLSTransform (he2024hlstransform, ) uses HLS to design a FPGA accelerator and synthesis combined with pipelining, memory unrolling, and memory partitioning and transfer optimizations, with the addition of 8-bit integer quantization (W8A8). On a tiny model with 110 million parameters, HLSTransform achieves 57.11 tokens/s on Xilinx Virtex UltraScale+ VU9P FPGA. SECDA-LLM (haris2024designing, ) utilizes quantization (W3A8) and designs an efficient FPGA-based LLM accelerators for the llama.cpp inference framework. By deploying on the PYNQ-Z1 board, it achieves 0.588 tokens/s for the TinyLlama model (1.1B). Chen et al. (chen2024understanding, ) investigate the feasibility and potential of model-speciic spatial acceleration for LLM inference on FPGAs. They introduce a comprehensive analytical model to estimate the LLM inference performance of FPGA accelerator with W4A8 quantization, and provide a library of high-level synthesis (HLS) kernels that are composable and reusable. For Llama2-7B, during prefilling phase, they can achieves about 213 tokens/s, 43 tokens/s, and 320 tokens/s on Xilinx Alveo U280, VCK5000, and VHK158 FPGAs, respectively. During decode stage, they can achieves about 200 tokens/s, 40 tokens/s, and 333 tokens/s on Xilinx Alveo U280, VCK5000, and VHK158 FPGAs, respectively.
3.1.5. ASIC
@@
Weight-Only Quantization. Despite the memory footprint reduction achieved by weight-only quantization, the actual computing performance is not really improved due to dequantization from integer to float. FIGNA (jang2024figna, ) proposes dedicated FP-INT arithmetic units designed specifically for FP-INT MAC operations and integrates them on the accelerator. FIGNA with FP16-INT4 provides 3.2768 TOPS computing power and 26.58W power consumption by considering all memory access in 28nm at 100MHz. Estimated result shows that for OPT-6.7B it can achieve 21.332 tokens/s in decode stage. Different from normal quantization methods, MECLA (qin2024mecla, ) proposes a parameter-efficient scaling sub-matrix partition method (SSMP) to decompose large weight matrices into several tiny-scale source sub-matrices (SS) and derived sub-matrices (DS). For memory issues, SSMP avoids accessing the full weight matrix but only requires small SS and DS scaling scalars. For computation issues, the proposed accelerator fully exploits the intermediate data reuse of matrix multiplication via on-chip matrix regrouping, inner-product multiplication re-association, and outer-product partial sum reuse. Totally, it can reduce 83.6% memory access and 72.2% computation. MECLA provides 14.008 TOPS computing power and $\sim$96W (1.9763W+94W) power consumption by considering all memory access under 28nm. For Llama2-7B and BLOOM-7B, compared to NVIDIA V100 GPU, MECLA achieves 6.74$\times$ and 5.91$\times$ inference speedup ($\sim$161 tokens/s and 141 tokens/s, respectively).
@@
Weight-Activation Quantization. Based on the key insight that outliers are important while the normal values next to them are not, OliVe (guo2023olive, ) adopts an outlier-victim pair (OVP) quantization and handles outlier values locally with low hardware overheads. This enables a memory-aligned W4A4/W8A8 quantization, which can be efficiently integrated to the existing hardware accelerators like systolic array and tensor core. OliVe provides 0.71 TOPS computing power and $\sim$8W (0.2806W+7.9872W) power consumption by considering all memory access under 22nm. Estimated results shows that for OPT-6.7B it can achieve 9.173 tokens/s in decode stage. Li et al. (li2024quantization, ) uniformly group weights and activations to ensure workload balance for hardware, and propose two approaches called channel sorting and channel selection to enhance the performance of quantization. It provides 1.43 TOPS computing power and $\sim$8.5W (0.472W+7.9872W) power consumption by considering all memory access under 65nm. Estimated results shows that for OPT-6.7B it can achieve 19.733 tokens/s in decode stage. Tender (10609625, ) decomposes weight and activation matrices by groups with different size to smooth the impact of outliers. And the format of scale factors are powers of two apart, which avoids explicit dequantization and extension to the commodity tensor compute hardware. It is 7.174W (1.60W+5.574W) power consumption by considering HBM2 memory access under 28nm. Result shows that for OPT-6.7B, Tender achieves 1.33$\times$ speedup (53.33 tokens/s) than NVIDIA A100 GPU.
3.1.6. PIM/NDP
@@
Figure 4. LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with quantization.
3.1.7. Quantitative Comparison
We first compare the power consumption, inference speed and energy efficiency for different hardware platforms, in Figure 4. For quantization, power consumption ranges from 3W to 450W, with inference speeds between 3 tokens/s and 1998 tokens/s. The energy efficiency ranges from 0.0167 tokens/J to 46.66 tokens/J. T-MAC (wei2024t, ) (CPU) achieves the lowest power consumption with 3W and Guo et al. (guo2024towards, ) (PIM/NDP) achieves the highest throughput with batch size 1 (pre-silicon simulation result). And Guo et al. (guo2024towards, ) also achieves the highest energy efficiency with 46.66 tokens/J.
@@ 3.2.1. Overview
Figure 5. Sparsity and sparsity patterns.
Table 4. Sparsity on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@ 3.2.2. CPU
Activation Sparsity. Activation sparsity is determined by activation functions. Commonly using SwiGLU (shazeer2020glu, ) and GeGLU (hendrycks2016gaussian, ) exhibits limited sparsity for LLMs, but simply replacing these functions with ReLU fails to achieve sufficient sparsity. Turbo Sparse (song2024turbo, ) proposes the dReLU activation function to improve LLM activation sparsity, along with a high-quality training data mixture ratio to facilitate effective sparsity. By applying their sparsity method to the Mistral and Mixtral models, only 2.5 billion (35.7%) and 4.3 billion (9.2%) parameters are activated per inference iteration, respectively. For Mistral-7B, Turbo Sparse achieves 8.71 tokens/s and 9.94 tokens/s on Intel i9-14900HX processor and Intel i7-12700K processor, respectively. For Mixtral-47B with 4-bit quantization, Turbo Sparse achieves 16.1 tokens/s, 11.98 tokens/s and 11.1 tokens/s on Intel i9-14900HX, Intel i7-12700K and SnapDragon 8 Gen3, respectively. ProSparse (song2024prosparse, ) also introduces activation function substitution, progressive sparsity regularization, and activation threshold shifting to help non-ReLU LLMs obtain high activation sparsity without performance degradation. For Llama2-7B and Llama2-13B, ProSparse achieves high sparsity of 89.32% and 88.80%, and 16.3 tokens/s and 8.67 tokens/s, respectively, based on PowerInfer (song2023powerinfer, ) framework.
3.2.3. GPU
@@
Weight Sparsity. LLM-pruner (ma2023llm, ), adopts structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the majority of the LLM’s functionality. To this end, the performance of pruned models can be efficiently recovered through tuning techniques, LoRA, in merely 3 hours, requiring only 50K data. We validate the LLM-Pruner on three LLMs, including Llama, Vicuna, and ChatGLM, and demonstrate that the compressed models still exhibit satisfactory capabilities in zero-shot classification and generation.
@@
LLMs can be pruned to at least 50% sparsity in one-shot, without any retraining, at minimal loss of accuracy. SparseGPT (frantar2023sparsegpt, ) requires a sophisticated weight update procedure in an iterative pruning process. Wanda (sun2023wanda, ) prunes weights with the smallest magnitudes multiplied by the corresponding input activations, on a per-output basis. Notably, Wanda requires no retraining or weight update, where pruning process is faster. Besides unstructured pattern, these two methods generalizes to semi-structured N:M (2:4 and 4:8) patterns. E-Sparse (li2023sparse, ) introduces entropy to quantify the information richness within each channel (intra-channel) of the input features, and adopts it to enhance the feature norms (crosschannel) as a metric to evaluate parameter importance. Furthermore, it proposes Channel Shuffle to reorder the information distribution in LLMs to obtain N:M Sparsity with less information loss. 2:4 sparsity as supported by NVIDIA GPUs of generation Ampere and newer theoretically offers 2$\times$ acceleration of matrix multiplications. In practical, 2:4 sparsity can achieve 1.54$\times$-1.79$\times$ speedup for MatMul, and end-to-end speedups are about 1.21$\times$-1.25$\times$ (due to some extra overheads from e.g. attention).
@@
Based on the key observation that the bottleneck of LLM inference is the skinny matrix multiplications, Flash-LLM (xia2023flash, ) proposes a general Load-as-Sparse and Compute-as-Dense methodology for unstructured sparse matrix multiplication. Flash-LLM proposes a new sparse format called Tiled-CSL to relieve the memory bandwidth bottleneck and support the tile-by-tile SpMM execution with tensor cores. For OPT-30B, Flash-LLM achieves 80% sparsity with 1.44% accuracy decrease and about 290 tokens/s, 500 tokens/s, 800 tokens/s, and 1187 tokens/s on single A100 GPU with batch sizes 8, 16, 32, and 64, respectively.
@@
Agarwalla et al. (agarwalla2024enabling, ) combine the SparseGPT one-shot pruning method and sparse pretraining to pretrain a high sparsity LLM. They deploy model on GPU and CPU by utilizing Neural Magic’s DeepSparse engine and Neural Magic’s nm-vllm engine, respectively. For Llama-7B, on NVIDIA A10 GPU, they achieve 44.4 tokens/s and 47.9 tokens/s with 50% sparsity and 70% sparsity, respectively. On AMD EPYC 9R14 Processor, they achieve 4.4 tokens/s and 6.9 tokens/s with 50% sparsity and 70% sparsity, respectively.
@@
Existing methods require costly retraining, forgo LLM’s in-context learning ability, or do not yield wall-clock time speedup on modern hardware. DejaVu (liu2023deja, ) predicts contextual sparsity on the fly given inputs to each layer, along with an asynchronous and hardware-aware implementation that speeds up LLM inference. For OPT-175B model, DejaVu achieves up to 75% sparsity and 50 tokens/s on 8 A100-80GB GPUs with batch size 1, which is over 2$\times$ and 6$\times$ faster than FasterTransformer and Hugging Face implementation, respectively.
@@
Attention Sparsity. During the prefilling phase of LLM inference, attention computation complexity scales quadratically with input sequence length. Given limited GPU computing and memory resources, attention sparsification can reduce the number of attention values to accelerate prefilling phase. For static sparsity, Sparse Transformer (child2019generating, ), StreamingLLM (xiao2023streamingllm, ), Bigbird (zaheer2020big, ), and Longformer (beltagy2020longformer, ) use the manual combination of global and local patterns to replace the full attention patterns. The local pattern captures the local context of each token within a fixed size or stride while the global pattern captures the relationship between the specific tokens to all other tokens. For Llama2-7B and Llama2-13B models, StreamingLLM achieves 15.38-32.26 tokens/s and 9.43-20.83 tokens/s on single NVIDIA A6000 GPU, respectively. For dynaimic sparsity, Adaptively Sparse Attention (correia2019adaptively, ) replaces softmax with $\alpha$-entmax, a differentiable generalization of softmax that allows low-scoring words to receive precisely zero weight and drops parts of the context that are no longer required for future generation. Reformer (kitaev2020reformer, ) replaces dot-product attention by using locality-sensitive hashing, changing the complexity from $O(L^{2})$ to $O(LlogL)$, where $L$ is the sequence length. Sparse Flash Attention (pagliardini2023faster, ) extends FlashAttention (dao2022flashattention, ) GPU kernel and encompasses key/query dropping and hashing-based attention. Sparse Sinkhorn Attention (tay2020sparse, ) adopts a learned sorting network to align keys with their relevant query buckets, ensuring that attention is computed only between the corresponding query-key pairs. H2(zhang2024h2o, ) observes that a small portion of tokens (called Heavy Hitters, H2) contributes most of the value when computing attention scores. H2O introduces a dynamic attention sparsification method to adopt KV cache eviction policy that dynamically retains a balance of recent and H2 tokens. For OPT-6.7B model, H2O with 20% H2 achieves 30.4 tokens/s on single NVIDIA T4 GPU.
3.2.4. FPGA